OpenStudy (anonymous):

Of the following properties, distributive, commutative, or associative, match the example with the property it illustrates. A. 1+(3+4) = (1+3)+4 B. 5 + 3 = 3 + 5 C. 3(2+5)=3*2+3*5

OpenStudy (zzr0ck3r):

(a+b)+c = a+(b+c) associative a+b = b+a commutative a(b+c) = ab+ac distributive

OpenStudy (anonymous):

Thank you!!!